OpenStudy (anonymous):

A Ferris wheel with a diameter of 50 feet has 30 equally spaced passenger cars. What is the length of the arc between the centerline of two consecutive cars? Use the value π = 3.14, and round your answer to the nearest tenth of a foot.

OpenStudy (misty1212):

the circumference of the circle is \(2\pi r\) or in your case \(100\pi\)

OpenStudy (misty1212):

divide that by 30

OpenStudy (michele_laino):

If L is the requested length, then we can write: \[\large 2 \times 3.14 \times \frac{D}{2} = 30 \times L\] where D is the diameter of your wheel

OpenStudy (anonymous):

ohh thanks you guys

OpenStudy (michele_laino):

thank you!

OpenStudy (anonymous):

so the answer is a ? im confused whats the answer